Six sales representatives for a company are to be chosen at random to participate in a training program. The company has twelve sales​ representatives, two in each of six regions. What is the probability that the six sales representatives chosen to participate in the training program will be from only three of the six ​regions?


1
Expert's answer
2021-12-14T06:14:25-0500

"P=\frac{C^6_6}{C^6_{12}}=\frac{1}{C^6_{12}}=\frac{6!6!}{12!}=0.0011"